OpenStudy (anonymous):

write the first five terms of the arithmetic sequence whose nth term is given an=20n+100

OpenStudy (anonymous):

Put n = 1 for first term \[a_n = 20n + 100\] Put n = 1 here: \[a_1 = 20(1) + 100\] \[a_1 = 20 + 100\] \[\color{blue}{a_1 = 120}\] Similarly by putting n = 2, 3, 4 and 5, find the 2nd, 3rd, 4th and 5th term respectively...

OpenStudy (anonymous):

And more important thing: \[\huge \color{green}{\textbf{Welcome To Openstudy...}}\]

OpenStudy (anonymous):

Thank you Waterineyes. That was the same answer I got. I wanted to see if I was correct or not.

OpenStudy (anonymous):

Then what you got as your 2nd, 3rd, 4th and 5th terms??

OpenStudy (anonymous):

I got 120, 140, 160, 180, and 200

OpenStudy (anonymous):

Good... Well Done.. Keep it up..